# Even more efficient magic state distillation by zero-level distillation

### v1

@@ IV Direct Code Conversion from Steane to surface codes
Here we present another protocol to prepare the logical magic state on the rotated surface code by converting the Steane code state to the rotated surface code directly without teleportation. This approach does not require the generation of two logical states and hence consumes fewer qubits compared to the teleportation-based approach above. As a drawback, the circuit depth is increased to 42, which is approximately twice as long as the teleportation-based approach. As a result, the logical error rate is worse than that of the teleportation method due to the strong influence of idling noise.
@@
First, as shown in Fig. 9(LABEL:sub@fig:c1), a magic state is encoded as described in Sec III. After that, as shown in Fig. 9 (LABEL:sub@fig:c2)–(LABEL:sub@fig:c12), we perform one-bit teleportations to move the qubits appropriately and stabilizer measurements to detect errors, and to form the stabilizer operators for the rotated surface codes. Specifically, Fig. 9(LABEL:sub@fig:c3)–(LABEL:sub@fig:c5) show $Z$, $X$, $Z$-stabilizer measurements respectively. If the syndrome values are odd in these three stabilizer measurements, the distillation process must be restarted from the beginning.
@@
Next, we deform the Steane code to a 6-qubit code of distance two while keeping the logical state and fault-tolerance. As shown in As shown in Fig. 9(LABEL:sub@fig:c6), qubit 6 is measured directly in the $X$-basis. The 6-qubits code consists of two $X$ and $Z$ weight-four stabilizers and one weight-three $X$ stabilizer originating from the Steane code as shown in Fig. 9(LABEL:sub@fig:c7). Similar to the Steane code, the logical $Z$ operator is composed of three qubits, $Z_{1}Z_{3}Z_{5}$. On the other hand, the logical $X$ operator is composed of two qubits, $X_{1}X_{4}$. In Fig. 9(LABEL:sub@fig:c8), certain qubits are prepared in the $\ket{+}$ states, which are used to form an $X$ stabilizer of the rotated surface code later.
@@
Finally, the 6-qubit code is deformed to the rotated surface code by measuring appropriate stabilizers with obtaining the syndrome values inherited from the Steane code as follows. In Fig. 9(LABEL:sub@fig:c9), two $Z$-type operators are measured; one is for the stabilizer for the rotated surface code and the other is to convert the $X$-type stabilizer of the Steane code (red face) into $X$ stabilizers of the rotated surface code. The measurement outcomes of these two $Z$-type operators have to coincide since they are linked by the blue and green stabilizers of the Steane code. Therefore, if the measurement outcomes disagree indicating any error, and hence the distillation process is rejected. Similarly, in Fig. 9(LABEL:sub@fig:c10), two $X$-type operators are measured to form the $X$ stabilizers for the rotated surface code. If their measurement outcomes do not coincide, the distillation process is rejected. This measurement removes the $Z$ stabilizer on the blue face, and only the $X$ stabilizer is left (colored yellow). In Fig. 9(LABEL:sub@fig:c11), the $X$-stabilizers are measured to be formed, which is repeated twice to detect measurement errors. In Fig. 9(LABEL:sub@fig:c12) two $Z$-stabilizer stabilizers are measured and formed, which are repeated twice to detect measurement errors. This measurement removes the $X$ stabilizer on the green face, and only the $Z$ stabilizer is left (colored orange). This completes the code conversion and the magic state is now encoded in the rotated surface code. 42 physical depths are used to complete the code conversion including the magic state distillation part on the Steane code. Throughout this process, the Pauli frame should be updated based on the measurement outcomes as usual.
@@ V Numerical simulation
We perform numerical simulation to verify zero-level distillation works fault-tolerantly and to estimate the logical error rate and success probability. Specifically, since our protocol includes non-Clifford gates, we employ a full vector simulation using Qulacs [25]. While the original circuits require about 40 qubits for the rotated surface code (50 for the planar surface code), we perform numerical simulations using only about 20 qubits by reusing the ancilla qubits without changing the structure of the original circuit [26]. In the case of the code conversion, 15 qubits are enough to simulate.
@@
Regarding the noise model, each single-qubit and two-qubit gates are followed by single-qubit and two-qubit depolarizing noise with a probability $p$, respectively, where an idling process is also regarded as an identity gate and hence is followed by noise. In order to estimate the logical error rate $p_{L}$, the obtained magic state on the rotated surface code is virtually projected to the code space, and fidelity between the ideal magic state is calculated. If fidelity is not a unit, we count it as a logical error. The reason why we project the state to the ideal code space is to estimate the logical error rate by excluding potentially detectable errors. These detectable errors can be post-selected in the further process in the leading order by sacrificing a small amount of success probability.
@@
The logical error rate $p_{L}$ is estimated for various physical error rates $p$, where the numbers of samples are chosen to be $10^{6}$ to $10^{7}$. The resultant logical error rate is shown in Fig. 10 for the cases of teleportation-based approaches for the planar (red) and rotated (blue) surface codes, and the code conversion approach (purple). We can see that the logical error rate scales like $p_{L}=a\times p^{2}$ and the coefficient is $93.4$ for the planar surface code and 106 for the rotated surface code. In the case of the code conversion, the coefficient is given by $199$. Compared to the rotated surface code, the planar surface code has a slightly lower logical error rate due to less shuttling of qubits. The intersection of $p_{L}\simeq 100p^{2}$ and $p_{L}=p$ is located around $p=10^{-2}$, which indicates that the logical error rate is improved by one order of magnitude at the times of $p=10^{-3}$ and by two orders of magnitude at the times of $p=10^{-4}$. Compared to the teleportation-based approach, the code conversion provides a slightly higher logical error rate because of its large circuit depth.
@@
The success probabilities of distillation are plotted as functions of the physical error rate $p$ in Fig. 11 for the teleportation-based approaches of the planar surface code (red) and the rotated surface code (blue), and the code conversion (purple). The success rate is 70% when $p=10^{-3}$ and 95% when $p=10^{-4}$, indicating that distillation succeeds with a high probability. Compared to the planar surface code, the rotated surface code has a slightly higher success probability since it uses less number of physical qubits. The success probability of the code conversion is almost the same as that of the planar surface code since it uses fewer number of physical qubits and gates.
@@
Figure 10: The logical error rate $p_{L}$ is plotted as a function of the physical error rate $p$ for the teleportation-based approaches of the planar (red circle) and rotated (blue cross) surface code, and the code conversion (purple square). The red solid, blue dashed, and purple dotted lines are fitting with $ap^{2}$. The green dash-dotted line indicates $p_{L}=p$.
@@
Figure 11: The success rate is plotted as a function of the physical error rate $p$ for the teleportation-based approaches of the planar (red circle) and rotated (blue cross) surface code, and the code conversion (purple square).
VI conclusion and Discussion
@@
We proposed zero-level distillation, which efficiently distills and prepares logical magic state encoded on the surface codes without using multiple logical qubits. All operations required can be implemented on the square lattice connectivity and the number of required physical qubits is substantially reduced and spatial overhead for one or two logical patches is enough.
@@
According to the numerical simulation, zero-level distillation with teleportation successfully reduces the logical error rate $p_{L}$ of a logical magic state to $p_{L}=100\times p^{2}$. For example, when $p=10^{-3}$ and $p_{L}=10^{-4}$, the logical error rates result in $p=10^{-4}$ and $p_{L}=10^{-6}$, respectively indicating one and two orders of magnitude improvement. The success probability is reasonably high even when $p=10^{-3}$. The depth of the zero-level distillation circuit is only 25, and hence it is compatible with the conventional multi-level distillation routines [20]. In addition, we also developed zero-level distillation based on the code conversion to further reduce the number of physical qubits employed.
Let us discuss how the proposed zero-level distillation has an impact when it is combined with further multi-level distillation. By changing the rotation axis, we can distill |$\ket{T}$ magic states with zero-level distillation. Such magic states can be used as level-1 magic states of multi-level distillation protocols generating |$\ket{T}$ magic states. For example, with $p=10^{-4}$, zero-level distillation reduces the spatiotemporal overhead of two-level 15-to-1 distillation protocols in Ref. [20] achieving $p_{L}=1.5\times 10^{-15}$ by more than 50%. A detailed analysis should be presented in future work. In addition to the non-Clifford gates, by using the transversality of $S$ gate in the Steane code, we can also apply our protocol for preparing logical $Y$ eigenstate reliably on the surface code, which could reduce the overhead for the logical $S$ gate for the surface codes.
@@ Acknowledgements.
This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118067394 and JPMXS0120319794, JST COI-NEXT Grant No. JPMJPF2014, and JST Moonshot R&D Grant No. JPMJMS2061.
@@ Appendix A Zero-level distillation for the planar surface code
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Qubits arrangement and measurements during teleportation from the Steane code to the planar surface code. The white dots and numbers correspond to the locations of the qubits shown in Fig. 3, where the distilled magic state is encoded in the white qubits. The light blue arrows indicate the transfer of qubits using one-bit teleportations. The red, blue, and green faces indicate the stabilizers of the Steane code, and the orange and yellow faces indicate the stabilizers of the planar surface code. The red, blue, and green dots are the ancilla qubits for the stabilizer measurements of the Steane code. The purple faces indicate the $Z$-stabilizers at the boundary for the lattice surgery. The gray lines indicate logical operators of the Steane code. The purple box and the gray box indicate logical operators of the planar surface code.
The magic state distilled on the Steane code can also be teleported to the planar surface code as well as the rotated surface code. In this case, the |$\ket{+}_{L}$ state encoded with the planar surface code and $\ket{+}$ ancilla physical qubits are prepared in parallel with the Hadamard test as shown in Fig. 12(LABEL:sub@fig:planer_1). After the Hadamard test, some data qubits are moved as shown in Fig. 12(LABEL:sub@fig:planer_2). Then, as shown by the purple area of Fig. 12(LABEL:sub@fig:planer_3), $L_{Z}^{\mathrm{Steane}}\otimes L_{Z}^{\mathrm{Surface}}$ is measured as a lattice surgery. During this lattice surgery, the $Z$-stabilizers have to be measured twice to detect measurement errors. As shown by the purple dots in Fig. 12(LABEL:sub@fig:planer_4), ancilla qubits on the boundary are measured in the $X$ basis, which breaks the $Z$-stabilizers on the boundary, and two code states are split.
As shown in Fig. 12(LABEL:sub@fig:planer_4), the $Z$ stabilizers are measured on both Steane and planar surface codes. Finally, the qubits on the Steane code are directly measured in the $X$ basis to calculate the parities of $X$-stabilizers and logical operators. Throughout all operations from the beginning, 23 physical depths are used. Note that the Pauli frame has to be updated appropriately throughout these measurements as usual.

### v2

@@ IV DIRECT CODE CONVERSION FROM STEANE TO SURFACE CODES
We propose an alternative protocol to prepare the logical magic state on the rotated surface code by converting the Steane-code state to the rotated surface code directly, without relying on teleportation. This approach eliminates the need to generate two logical states and hence consumes fewer qubits compared to the teleportation-based approach described above. As a drawback, the circuit depth is increased to 42, roughly double that of the teleportation method. This leads to a higher logical error rate caused by additional idling noise.
@@
First, as shown in Fig. 10(a), the magic state is encoded as described in Sec III. Subsequently, as shown in Fig. 10 (b)–(l), we perform one-bit teleportations to move the qubits appropriately and stabilizer measurements to detect errors, and to form the stabilizer operators for the rotated surface code. Specifically, Fig. 10(c)–(e) show $Z$-, $X$-, and $Z$-stabilizer measurements, respectively. If any of the syndrome values is odd in these three stabilizer measurements, the distillation process must be restarted from the beginning.
@@
Next, we deform the Steane code to a six-qubit code of distance two while preserving the logical state and fault tolerance. As shown in Fig. 10(f), qubit 6 is measured directly in the $X$ basis. The six-qubit code consists of two weight-four $X$ and $Z$ stabilizers and one weight-three $X$ stabilizer originating from the Steane code as shown in Fig. 10(g). Similar to the Steane code, the logical $Z$ operator is composed of three qubits, $Z_{1}Z_{3}Z_{5}$. On the other hand, the logical $X$ operator is composed of two qubits, $X_{1}X_{4}$. In Fig. 10(h), certain qubits are prepared in the $\ket{+}$ states, which subsequently form part of an $X$-stabilizer in the rotated surface code.
@@
Finally, the six-qubit code is deformed to the rotated surface code by measuring appropriate stabilizers with obtaining the syndrome values inherited from the Steane code. In Fig. 10(i), two $Z$-type operators are measured; one is for the stabilizer of the rotated surface code and the other is to convert the $X$-type stabilizer of the Steane code (red face) into $X$ stabilizers of the rotated surface code. The measurement outcomes of these two $Z$-type operators have to coincide since they are linked by the blue and green stabilizers of the Steane code. Therefore, if the measurement outcomes disagree indicating any error, the distillation process is rejected. Similarly, in Fig. 10(j), two $X$-type operators are measured to form the $X$ stabilizers for the rotated surface code. If their measurement outcomes do not coincide, the distillation process is rejected. This measurement removes the $Z$ stabilizer on the blue face, and only the $X$ stabilizer is left (colored yellow). In Fig. 10(k), the $X$ stabilizers are measured to be formed, which is repeated twice to detect measurement errors. In Fig. 10(l) two $Z$-type stabilizers are measured and formed, which are repeated twice to detect measurement errors. This measurement removes the $X$ stabilizer on the green face, and only the $Z$ stabilizer is left (colored orange). This completes the code conversion, and the magic state is now encoded in the rotated surface code. A physical depth of 42 is used to complete the code conversion, including the magic state distillation part on the Steane code. Throughout this process, the Pauli frame should be updated based on the measurement outcomes as usual.
@@ V NUMERICAL SIMULATION
We perform numerical simulations to verify the fault tolerance of zero-level distillation and to estimate the resulting logical error rate and success rate. Since our protocol includes non-Clifford gates, we employ full-vector simulation using Qulacs [28]. While the original circuits require approximately 40 qubits for the rotated surface code (50 for the planar surface code), we perform numerical simulations using only approximately 20 qubits by reusing the ancilla qubits without changing the structure of the original circuit [29]. In the case of the code conversion, 15 qubits are enough for the entire simulation.
Regarding the noise model, we employ a standard depolarizing noise model; each single-qubit and two-qubit gate is followed by single-qubit and two-qubit depolarizing noise with a probability $p$, respectively, where an idling process is also regarded as an identity gate and hence is followed by noise. Note that while this noise model has been widely used to evaluate the performance of existing fault-tolerant schemes, there is another noise model that is more realistic for superconducting qubit hardware such as SI1000 [30]. In such a noise model, it would be better to implement the qubit transportation with a direct swap operation to avoid measurements as mentioned before. Also, we expect that applying SI1000 would lead to an improved logical error rate, since the zero-level distillation circuit includes quite a few idling operations, which have less impact with such an error model with much smaller idling noise.
@@
In order to estimate the logical error rate $p_{L}$, the obtained magic state on the rotated surface code is virtually projected to the code space, and fidelity between the obtained magic state and the ideal magic state is calculated. If the fidelity is not a unity, we count it as a logical error. The reason why we project the state to the ideal code space is to estimate the logical error rate by excluding potentially detectable errors. These detectable errors can be postselected in the further process in the leading order by sacrificing a small amount of success rate.
@@
The logical error rate $p_{L}$ is estimated for various physical error rates $p$, where the numbers of samples is chosen to be between $10^{6}$ and $10^{7}$. The resultant logical error rate is shown in Fig. 11 for the cases of teleportation-based approaches for the planar (red) and rotated (blue) surface codes, and the code-conversion approach (purple). We can see that the logical error rate scales as $p_{L}=a\times p^{2}$ and the coefficient is $93.4$ for the planar surface code and 106 for the rotated surface code. In the case of the code conversion, the coefficient is given by $199$. Compared to the rotated surface code, the planar surface code has a slightly lower logical error rate due to less shuttling of qubits. The intersection of $p_{L}\simeq 100p^{2}$ and $p_{L}=p$ is located around $p=10^{-2}$, which indicates that the logical error rate is improved by 1 order of magnitude at $p=10^{-3}$ and by 2 orders of magnitude at $p=10^{-4}$. Compared to the teleportation-based approach, the code conversion provides a slightly higher logical error rate because of its large circuit depth.
@@
The success rate of distillation is plotted as a function of the physical error rate $p$ in Fig. 12 for the teleportation-based approaches of the planar surface code (red) and the rotated surface code (blue), and the code conversion (purple). The success rate is 70% when $p=10^{-3}$ and 95% when $p=10^{-4}$, indicating that distillation succeeds with a high probability. Compared to the planar surface code, the rotated surface code has a slightly higher success rate since it uses fewer physical qubits. The success rate of the code conversion is almost the same as that of the planar surface code since it uses fewer physical qubits and gates.
@@
Figure 11: The logical error rate $p_{L}$ plotted as a function of the physical error rate $p$ for the teleportation-based approaches of the planar (red circle) and rotated (blue cross) surface code, and the code conversion (purple square). The red, blue, and purple lines scale quadratically. The green line indicates $p_{L}=p$.
@@
Figure 12: The success rate plotted as a function of the physical error rate $p$ for the teleportation-based approaches of the planar (red circle) and rotated (blue cross) surface code, and the code conversion (purple square).
We also perform numerical simulations to verify that surface-code expansion works fault tolerantly and to estimate the logical error rate and success rate in such a case. For this purpose, a full-vector simulation using Qulacs cannot be used because of the large number of qubits. Instead, we use the stabilizer simulator Stim [31]. We replace all $e^{-i\frac{\pi}{8}Y}$ gates with $e^{-i\frac{\pi}{4}Y}$ gates and perform $e^{-i\frac{\pi}{4}Y}$ distillation instead of $e^{-i\frac{\pi}{8}Y}$ distillation. Since this change does not affect the circuit configuration, it is reasonable to estimate the logical error rate. In order to verify this, we compare the results obtained by the full vector simulation with Qulacs and Clifford simulation with Stim for the $d=3$ case without the expansion. Fig. 13 shows the difference in the logical error rate between two simulation approaches for $e^{-i\frac{\pi}{4}Y}$ and $e^{-i\frac{\pi}{8}Y}$ distillations. We find that replacing $\pi/8$ with $\pi/4$ slightly decreases the logical error rate, but the success rate results in the same, as shown in Fig. 14.
Figure 13: The logical error rate $p_{L}$ plotted as a function of the physical error rate $p$ for the $e^{-i\frac{\pi}{8}Y}$ distillation in Qulacs (blue cross), the $e^{-i\frac{\pi}{4}Y}$ distillation in Qulacs (black square), and the $e^{-i\frac{\pi}{4}Y}$ distillation in Stim (cyan dot). The blue, black, and cyan lines scale quadratically.
Figure 14: The success rate plotted as a function of the physical error rate $p$ for the $e^{-i\frac{\pi}{8}Y}$ distillation in Qulacs (blue cross), the $e^{-i\frac{\pi}{4}Y}$ distillation in Qulacs (black square), and the $e^{-i\frac{\pi}{4}Y}$ distillation in Stim (cyan dot).
Fig. 15 shows the results of the logical error rate when the code distance of an output state is extended. The cyan dots indicate the logical error rate when the code distance is not extended, and other points (blue, purple, orange, red) indicate the data with the code distances extended from $d=3$ to $d=5$ or $d=7$. The cyan, blue, and purple points are calculated using error detection when projecting to the code space. On the other hand, the orange and red points are calculated using error correction. Fig. 16 shows the results of the success rate when the code distance of an output state is extended. As Fig. 15, five data points are plotted for each code distance and projection mode. These results clearly demonstrate that the proposed method maintains a low logical error rate while slightly reducing the success rate, even when the surface code is expanded to mitigate the accumulation of logical errors after zero-level distillation.
Figure 15: The logical error rate $p_{L}$ plotted as a function of the physical error rate $p$. Five data are plotted for each code distance and projection mode. The detection and correction modes mean performing error detection and correction, respectively when expanding the code. Cyan, blue, purple, orange, and red lines scale quadratically. The green line indicates $p_{L}=p$.
Figure 16: The success rate plotted as a function of the physical error rate $p$ for each code distance and projection mode. The detection and correction modes mean performing error detection and correction, respectively when expanding the code.
VI QUALITATIVE IMPLICATIONS OF ZERO-LEVEL DISTILLATION
The potential implications of zero-level distillation include both early-FTQC [23] and full-fledged FTQC. In early-FTQC, the limited availability of physical qubits prevents the execution of conventional multilevel distillation protocols [24]. Therefore, zero-level distillation, which operates with a very small number of qubits, spatial overhead of almost one logical qubit, is well suited for early-FTQC, despite its worse scaling $100p^{2}$ compared to $35p^{3}$ of conventional methods. Zero-level distillation can generate magic states with a logical error rate of $10^{-6}$ under the physical error-rate assumption of $10^{-4}$, which would be sufficient to achieve Megaquop [32] in early FTQC. When applying continuous rotation gates with an approximation accuracy of $10^{-6}$, roughly speaking, a few tens of $T$ gates are required [33]. By employing zero-level distillation, it becomes possible to perform around a few $10^{4}$ continuous rotation gate operations with fully protected Clifford gates, thereby expanding the range of algorithms beyond NISQ.
Regarding the scalability to larger code distances for full-fledged FTQC applications, two studies, (0+1)-level distillation [34] and magic state cultivation [35], have already been discussed since the first version of this paper appeared. (0+1)-level distillation utilizes the magic states generated by zero-level distillation as input states for conventional level-1 distillation protocols. Combining zero-level distillation, whose logical error rate scales as $O(p^{2})$, with conventional methods scaling as $O(p^{3})$ results in an overall scaling of $O(p^{6})$. Zero-level distillation requires almost no additional qubits compared to conventional methods, resulting in a minimal increase in spatial overhead. Notably, Ref. [34] has demonstrated an improvement in the logical error rate by up to 6 orders of magnitude using this method, while consuming the same spatiotemporal overhead $2.5\times 10^{5}$ qubitcycles. Additionally, it is shown that the spatiotemporal overhead required to achieve a logical error rate of $10^{-16}$ can be reduced to approximately one third [34].
On the other hand, magic state cultivation is inspired by the concept of zero-level distillation as mentioned in Ref. [35]. It employs larger color codes to perform physical-level distillation even under connectivity constraints. This method successfully suppresses the logical error rate to $O(p^{5})$ by introducing various innovative techniques, such as the “double-check” method. As a result, it can generate magic states with a given logical error rate while reducing the spacetime overhead by 2 orders of magnitude [35].
VII CONCLUSION
@@
We proposed zero-level distillation, which efficiently distills and prepares the logical magic state encoded in surface codes without requiring multiple logical qubits. All operations required can be implemented on the square lattice connectivity and the number of required physical qubits is substantially reduced and spatial overhead for one or two logical patches is sufficient.
@@
According to the numerical simulation, zero-level distillation with teleportation successfully reduces the logical error rate $p_{L}$ of a logical magic state to $p_{L}=100\times p^{2}$. For example, when $p=10^{-3}$ and $p_{L}=10^{-4}$, the logical error rates result in $p=10^{-4}$ and $p_{L}=10^{-6}$, respectively, indicating 1 and 2 orders of magnitude improvement. The success rate is reasonably high even when $p=10^{-3}$. The depth of the zero-level distillation circuit is only 25, hence it is compatible with the conventional multilevel distillation routines [24]. In addition, we also developed zero-level distillation based on the code conversion to further reduce the number of physical qubits employed.
@@ Acknowledgements.
This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant JPMXS0120319794, JST COI-NEXT Grant No. JPMJPF2014, and JST Moonshot R&D Grant No. JPMJMS2061.
@@ Appendix A ZERO-LEVEL DISTILLATION FOR THE PLANAR SURFACE CODE
Figure 17: Qubits arrangement and measurements during teleportation from the Steane code to the planar surface code. The white dots and numbers correspond to the locations of the qubits shown in Fig. 3, where the distilled magic state is encoded in the white qubits. The light blue arrows indicate the transfer of qubits using one-bit teleportations. The red, blue, and green faces indicate the stabilizers of the Steane code, and the orange and yellow faces indicate the stabilizers of the planar surface code. The red, blue, and green dots are the ancilla qubits for the stabilizer measurements of the Steane code. The purple faces indicate the $Z$ stabilizers at the boundary for the lattice surgery. The gray lines indicate logical operators of the Steane code. The purple box and the gray box indicate logical operators of the planar surface code.
The magic state distilled on the Steane code can also be teleported to the planar surface code as well as the rotated surface code. In this case, the |$\ket{+}_{L}$ state encoded with the planar surface code and $\ket{+}$ ancilla physical qubits are prepared in parallel with the Hadamard test as shown in Fig. 17(a). After the Hadamard test, some data qubits are moved as shown in Fig. 17(b). Then, as shown by the purple area of Fig. 17(c), $L_{Z}^{\mathrm{Steane}}\otimes L_{Z}^{\mathrm{Surface}}$ is measured as a lattice surgery. During this lattice surgery, the $Z$ stabilizers have to be measured twice to detect measurement errors. As shown by the purple dots in Fig. 17(d), ancilla qubits on the boundary are measured in the $X$ basis, which breaks the $Z$ stabilizers on the boundary, and two code states are split.
As shown in Fig. 17(d), the $Z$ stabilizers are measured on both Steane and planar surface codes. Finally, the qubits on the Steane code are directly measured in the $X$ basis to calculate the parities of $X$ stabilizers and logical operators. Throughout all operations from the beginning, a physical depth of 23 is used. Note that the Pauli frame has to be updated appropriately throughout these measurements as usual.